OpenStudy (anonymous):

What is the value of x? x+5= -6x-16

OpenStudy (anonymous):

Explain

OpenStudy (igreen):

First we subtract 5 to both sides. \(x + 5 = -6x - 16\) \(x = -6x - 16 - 5\)

OpenStudy (anonymous):

so -6x - 16?

OpenStudy (igreen):

Subtract -16 - 5.

OpenStudy (anonymous):

then when i get that i subtract that answer by 5?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

-16-5=-21

OpenStudy (anonymous):

next i divide right?

OpenStudy (anonymous):

6x=-21

OpenStudy (anonymous):

then i divide -21 by 6 and i get -3.5?

OpenStudy (anonymous):

so my answer is -3.5?

OpenStudy (igreen):

Yes, -21 is correct..that gives us: x = -6x - 21 Now we add 6x to both sides, what do you get?

OpenStudy (anonymous):

-15

OpenStudy (anonymous):

correct or no?

OpenStudy (igreen):

No, we're adding 6x + x

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

you lost me there?

OpenStudy (igreen):

\(6x + x \rightarrow 6x + 1x \rightarrow 6 + 1\) Then we add the 'x' on the end..

OpenStudy (anonymous):

so we add 6+1 and we get 7

OpenStudy (igreen):

Yep, so 6x + x is 7x. So we have: \(7x = -21\) Now divide 7 to both sides.

OpenStudy (anonymous):

-3?

OpenStudy (igreen):

Yep, you got it. x = -3

OpenStudy (anonymous):

thanks

OpenStudy (igreen):

No problem, I have to go now..

OpenStudy (igreen):

Sorry.. :c

OpenStudy (anonymous):

aww

OpenStudy (igreen):

Bye!